OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

what is the slope you need to use to make it parallel to y=4x-2?

OpenStudy (anonymous):

It doesn't say..

OpenStudy (anonymous):

no you have the info you need, for it to be parallel you use the same slope as the equation given. so what is the slope in y=4x-2?

OpenStudy (anonymous):

would it be 4 ?

OpenStudy (anonymous):

hint in y=mx+b, m is the slope

OpenStudy (anonymous):

so then it would be 4??

OpenStudy (anonymous):

yes good job, so now we have y=4x+b, lets fill in the x and y from the coordinates given.

OpenStudy (anonymous):

so then 6=4(-3)+b?

OpenStudy (anonymous):

wait do you need to find the b also? do you have to graph it?

OpenStudy (anonymous):

No, but these are the choices they gave me to choose from. x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

no no you are on the right track! now solve for b in 6=4(-3)+b

OpenStudy (anonymous):

would b be -6?

OpenStudy (anonymous):